OpenStudy (anonymous):

Solve. Solve a + 1 = √b + 1 for b a) b = a2 + 2a b) b = a + 2a2 c) b = 2a2 + a

OpenStudy (anonymous):

√b + 1 = a + 1 √b = a b = a^2

OpenStudy (anonymous):

b=a2 + 2a?

OpenStudy (anonymous):

no you solve for b so you first need to get √b isolated so you subtract 1 from both sides getting a = √b and then squaring for b you get b = a^2